# Conceptualizing Trustworthiness and Trust in Communications

### v1

@@ Abstract
Trustworthiness and trust are basic factors in common societies that allow us to interact and enjoy being in crowds without fear. As robotic devices start percolating into our daily lives they must behave as fully trustworthy objects, such that humans accept them just as we trust interacting with other people in our daily lives.
How can we learn from system models and findings from social sciences and how can such learnings be translated into requirements for future technical solutions? We present a novel holistic approach on how to tackle trustworthiness systematically in the context of communications. We propose a first attempt to incorporate objective system properties and subjective beliefs to establish trustworthiness-based trust, in particular in the context of the future Tactile Internet connecting robotic devices. A particular focus is on the underlying communications technology.
@@ I Introduction — A User Perspective
We are in a phase where societies rely more and more on highly connected digital technologies as the key infrastructure for economic, education, and social functioning. However, only very few of us have the principal expertise to judge if these technologies are truly reliable in a very literal meaning. Even experts are typically unlikely to get access to the necessary and detailed information about a product or a system. This expertise and knowledge gap is a common dilemma we humans must deal with in our daily lives. The answer we have to this dilemma is trust. Trust is a basic ingredient for societies to function. So, how can we ensure that highly connected digital technologies have basic enablers in place to provide for such trust decisions by humans?
In this paper, we try to answer this question by proposing a model which incorporates the concepts of trust and trustworthiness. Our model is backed up by findings from social science as well as engineering research.
Highly connected digital technologies like remotely controlled vehicles will enable products with features that depend on connectivity, i.e. on the underlying communications system. Consequently, also the functional safety of such products will depend on the underlying communications system. However, once such technologies and products become key infrastructure, even social functioning, including social peace and freedom, are at stake and depend on the underlying communications systems.
5G cellular systems have enabled the Tactile Internet [4] in industrial and professional applications. With the coming of 6G cellular, we expect the Tactile Internet to move to consumers. As engineers and computer scientists, we know that successful attacks one these systems are realistic from a technology perspective.
@@
For over a century, humans have dreamed of personal robotic assistants that would make our lives easier. These robots could take on a variety of tasks, from cleaning our rooms to providing companionship. Hardly any of us can judge whether these robots are “good guys” or “bad guys” in a future digitization-shaped society. We call these future societies digitization-shaped since the foreseeable proliferation of digital products and services as “members” of society will be formative, and we humans will have to trust those digital “members” of society. However, rather than blindly trusting them we should base our trust decision on trustworthiness, i.e. objectively measurable properties of the system. But when are such systems trustworthy? 6G-enabled systems, as any technical system, can only “behave” human-like and hence, can only be trustworthy in a human (and humane) sense if engineers design them to be so. Engineers must make them trustworthy by design.
@@
It is not farfetched to assume that with the coming of 6G-cellular-enabled products, we humans will continue our practice of delegating our trust decisions and put our trust in certain organizations that place a label of trustworthiness on these new products. We will trust those organizations to have the technical ability to judge the very features of a product or service that ultimately are a proxy for their “behaviour” in human-machine interaction and ultimately in a future society.
Multiple questions arize thereof that we are tackling in the next sections, such as: What causes humans to put trust in each other or organizations? What influences our trust decisions? What are the building blocks of trustworthiness? We first look at the findings of social science regarding answers to these questions. Afterwards, we will adapt and extend the findings and transfer them to a technical context to answer the question of what we mean by a trustworthy “behaviour” of a product or service. To this end, we will propose a model which sets the concepts of “trust” and “trustworthiness” in relation.
@@
II A Peek at A Social Sciences Model of Trust
Figure 1: A model of trustworthiness and trust by Mayer et al. [9]
@@ II A Peek at A Social Sciences Model of Trust
Over the last few decades, social sciences and humanities have produced a multitude of definitions and explanations as well as empirical findings on the topic of trust [6]. Hence we reflect the experiences of humanities and social sciences in the area of trust and trustworthiness before conceptualizing them for the area of communications engineering.
@@
On a micro level, both psychology and educational science predominantly explore interpersonal trust in specific interaction situations. However, this micro perspective is not sufficient to convey the multiple facets of trust. A macro-level approach takes place primarily in sociology and political science, with a particular focus on system trust. Political science studies the concept of trust mainly from a developmental dynamic perspective, aiming to explain the development of trust in democracy or institutions. Economic science studies the concept of trust both from a basic theory and a problem-oriented perspective, the latter primarily in organizational and marketing research. Specifically, organizational research focuses on questions of what organizational trust is, what factors influence it, and what effects it can have on the organization. By contrast, philosophy investigates the connection between trust and morality as well as the relevance of trust for society. In comparison, communication science examines less the question of trust and more the credibility of the media.
@@
The spectrum of social science research on trust is very broad. Still, there has been a lack of consolidation of the research achievements of these different social science disciplines on the topic of trust. Therefore, and certainly owing to the different scientific perspectives and their varying emphases, there is no uniform understanding of the term of trust. Yet, the diverse disciplinary interest shows that trust is not a simple concept, but a heterogeneous and complex subject.
@@
To investigate the concept of trust, we use the widely received integrative trust model of Mayer et al. [9] on organizational trust. The focus of this model (Fig. 1), which analyses the circumstances of trust emergence, is primarily on the trustworthiness of the trusted party as perceived by the trust giver. The authors define trust as “the willingness of a party to be vulnerable to the actions of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party” [9]. Particular emphasis is placed on the moment of vulnerability and the willingness to take a certain risk. The integrative model explains the emergence of trust in a situation in which a relationship between the two parties has not yet evolved. The factors of perceived trustworthiness influence the emergence of trust. However, the expectation of trust does not yet involve any risk, because, according to the model, risk is only the result of trust. This view is in contradiction to the views of other authors, who presume a risk perception that requires trust in the first place.
@@
According to the model of Mayer et al. [9], the assessment of trustworthiness as the basis for an act of trust is influenced by the evaluation of the following three factors: ability, integrity, and benevolence. Abilities include skills, competencies, and attributes that enable a person to solve a specific problem. Benevolence describes how much the trusted party considers the well-being of the trusting party in his or her actions. Integrity implies qualities such as honesty and reliability. Although the factors can be considered separately, they are interrelated and can influence each other.
@@
In this model, trustworthiness is understood as a continuum. At the beginning of a relationship, the trust giver has no or very little information about the trust taker. In particular, the assessment of benevolence is difficult. Therefore, Mayer et al. assume that the perception of integrity is key to the initial trust building. As the relationship progresses, both actors obtain more information and gain a better assessment of ability and benevolence. Thus, as the relationship progresses positively, the degree of trust may increase.
@@
Parallel to the model of Mayer et al., the theory of public trust was developed in communication science, which also works with different trust factors [2]. A distinction is made between subject-specific trust factors (factual competence and problem-solving competence), socio-psychological trust factors (social behaviour, communication behaviour [transparency, consistency], character), and socio-normative trust factors (sense of responsibility, ethical behaviour). The more completely and intensively the trust factors apply, the more likely are trust gain, trust formation or trust constitution. If the factors apply only partially and/or to a lesser extent, the more likely it is to result in a reduction or loss of trust.
@@ III Trustworthiness and its Measurability from a technical perspective
If we shift the focus towards the interaction of humans with technical systems, the definition of trust as given in the previous section can basically remain the same, i.e. the willingness of a human to be vulnerable to the actions of a technical system based on the expectation that the system will perform a particular action important to the human, irrespective of the ability to completely monitor or control the technical system. But what about the trustworthiness of the technical system and its measurability?
@@ III-A Trustworthiness
The ISO/IEC standard TS 5723:2022 [8] describes trustworthiness as the “ability to meet stakeholders’ expectations in a verifiable way”. Similarly, NIST sees trustworthiness as an objective property of a system, since the “trustworthiness of a system is based on the concept of assurance” [10]. Note that other concepts like “quality” [7], “quality of service”, or “quality of experience” have a similar broad meaning according to their definitions. Nevertheless, it makes sense to have different terms, since the wording emphasizes different aspects. E.g., quality of service is often related to the fulfilment of non-functional requirements such as latency or bandwidth. While trustworthiness is often used interchangeably with security and privacy-related aspects, it is important to stress that the fulfilment of expectations of stakeholders of a given system spans a much broader scope of requirements for the design of a truly trustworthy system.
@@ III-B Characteristics of Trustworthiness and their measurability
However, with a broader scope, even more indicators should be included such as:
  • Benevolence (as used in social sciences)
@@
A TC is a specific characteristic of a system having specific objectives. Measuring a TC is done with the help of TC measurements while TC controls are used to implement a given TC.
The first challenge in reasoning about trustworthiness is to have a comprehensive list of TCs. The second challenge is to define the means of measurability of the TCs. For most of the TCs no generally accepted means for measurement exist. For some TCs (e.g. confidentiality or availability) the measurability appears to be more tangible than for others (like privacy or benevolence). In many cases, a combination of multiple measures is certainly necessary to obtain a comprehensive understanding of the trustworthiness at stake. The individual weighting of the specific TCs and the means of measurement certainly vary depending on the domain and the context in which trustworthiness is being evaluated.
@@ IV Measurability of Trustworthiness
Ideally, trustworthiness as a whole should be measurable and quantifiable to allow for comparing systems and to meet the requirement of verifiability of user expectations. Despite various proposals reported in the literature [3] to create a metric for trustworthiness, no convincing solution exists today. The challenges behind designing a metric for trustworthiness are manifold. One reason is that trustworthiness involves several different TCs – and many of them lack sound metrics. Take for instance privacy as one TC. Privacy itself comprizes different aspects such as anonymity and unlinkability. However, even for these more basic properties of privacy, besides isolated measures like $k$-anonymity or differential privacy, no universal applicable metrics exist [12]. Another challenge is related to the intended nature of a metric for trustworthiness. Should it be a single number allowing for easy comparison but limiting expressiveness? Should it be a vector containing a value for each TC? This would enhance the expressiveness but makes comparison more difficult. Taking e.g. consumer reports as an example we can often see a combination of both approaches: a single value (e.g. inspired by school marks) for easy and quick judgements, but additionally measurements for individual properties of the service or product in question allowing a consumer to adapt the “base judgement” according to their individual situation, beliefs, and requirements. Similar approaches can also be seen e.g. in the IT security domain. Here the Common Vulnerability Scoring System (CVSS) on the one hand provides a single score allowing system administrators to do a quick risk assessment, but on the other hand allows for individual adaptation based on individual requirements (e.g. which protection goals are important) and the individual situation (e.g. do countermeasures exist like firewalls, which mitigate exploitability). So, would it make sense to define a trustworthiness vector based on the measured TCs as well as give a weighted trustworthiness score?
@@
Another question regarding a sound trustworthiness metric is whether the trustworthiness value should be time-dependent or time-invariant. A time-dependent metric might be more intuitive in the first place since the judgement of whether e.g. a cryptographic algorithm can be considered trustworthy might change over time when new weaknesses of the algorithm are discovered (thanks to cryptanalysis) or thanks to new technological developments (e.g. quantum computers). However, a time-dependent metric would be rather useless if the related values would be valid only in the very moment they are measured (or calculated). Such a metric would especially not help guide humans regarding their trust decisions.
@@
Therefore, the trustworthiness metric needs to be designed in a way which leads to values which have a reasonable validity period. Consequently, a metric which leads to time-independent values (or at least to values which equally change over time for similar systems) would be desired. Taking a second look at the example of the cryptographic algorithm and considering the symmetric encryption algorithm DES as a specific one, one could argue, that the inherent security “level” provided by DES in terms of effort necessary to find the secret key (through calculations) did not change substantially since its introduction in the 1970s. What changes is the “perceived” or relative security given the technological advancements in computation power allowing (investing the same amount of money) to break DES today within minutes while in the 1970s it took years. Translating this to another domain: What was considered a “fast” car in the 1930s might be considered a “slow” car today – although the maximum speed of that car remains the same, i.e. is time-invariant.
@@
Given the observations above, a possible solution could be to have two kinds of trustworthiness metrics: one “absolute” trustworthiness value, which is time-invariant to reflect the essence of trustworthiness provided by the system and to allow for comparison, and a “relative” trustworthiness value which is time dependent but captures better the current level of trustworthiness to more easily support human trust decisions. Moreover, the need for the latter can even be derived from the definition of trustworthiness itself, which is the “ability to meet stakeholders’ expectations”. Therefore the “level of trustworthiness” can always be understood as relative to the expectation of a given user (stakeholder).
@@
The different TCs which influence the trustworthiness of a system are not independent of each other. This poses additional challenges on optimizing the overall trustworthiness of a system. Consider for example TCs which are directly related to protection goals from the IT security domain, e.g. confidentiality, integrity, accountability, availability, anonymity etc. There are certain dependencies among these protection goals [13], e.g. confidentiality and availability weaken each other. Therefore, the related TCs are also not independent from each other. But there exist even dependencies across different TCs: Given the current state of the art, security enhancements often have a negative impact on usability; increased privacy could lead to reduced functionality expressed by the “privacy vs. utility” trade-off during system design. Consequently, for an informed trust decision that we can rely on, the correlations of the TCs must be analysed and taken into account. It is important to understand specific definitions and measuring methods of the TCs to ensure the filtering and weighting of the TCs is carried out resulting in a meaningful trust decision that confirms the learnings and minimizes risks.

### v2

@@ Abstract
Trustworthiness and trust are fundamental factors in societies that enable us to interact and enjoy mingling in crowds without fear. As robotic devices start permeating our daily lives, they must behave as completely trustworthy objects so that people will accept them just as they would trust other people when interacting with them in their daily lives.
As trust and trustworthiness have been researched in social sciences for many years, this opens the question: How can we learn from system models and findings from social sciences to translate such learnings into requirements for future technical solutions? This is of particular importance now, as 5G and 6G cellular communications open the door for the Tactile Internet — connected robotics interacting with humans. We present a novel holistic approach on how to tackle trustworthiness systematically in the context of communications. We propose a first attempt to incorporate objective system properties and subjective beliefs to establish trustworthiness-based trust.
@@ I Introduction — A User Perspective
For over a century, humans have dreamed of personal robotic assistants that would make our lives easier. These robots could take on a variety of tasks, from cleaning our rooms to providing companionship. Hardly any of us can judge whether these robots are “good guys” or “bad guys” in a future digitization-shaped society. We call these future societies digitization-shaped since the foreseeable proliferation of digital products and services as “members” of society will be formative, and we humans will have to trust those digital “members” of society. However, rather than blindly trusting them, we should base our trust decisions on trustworthiness, i.e., objectively measurable properties of the system. But when are such systems trustworthy? 6G-enabled systems, as any technical system, can only “behave” human-like and hence, can only be trustworthy in a human (and humane) sense if engineers design them to be so. Engineers must build an understanding of this and make systems trustworthy by design.
@@
It is not far-fetched to assume that with the coming of 6G-cellular-enabled products, we humans will continue our practice of delegating our trust decisions, putting our trust in certain organizations that place a label of trustworthiness on these new products. We will trust those organizations to have the technical ability to assess the very features of a product or service that fundamentally are a proxy for their “behavior” in human-machine interaction and ultimately in a future society.
We are currently in an era where societies increasingly depend on interconnected digital communications as the backbone of our economy, education, government, as well as personal social interactions. However, hardly any of us have the expertise or the tools needed to assess whether our digital communicating world is genuinely reliable in the most literal sense. Even experts often lack access to the comprehensive and detailed information required to fully evaluate a complex product or system. Therefore, when a decision needs to be made, as humans, we simplify the complex problem of understanding all risk factors by trusting most and concentrating on understanding the risk of the crucial variables [12]. There are many examples in our daily life. We trust a car when getting in that it is safe; we trust to drive on one side and that other cars stick to the same side; this way we can focus our effort on driving the car safely. Trust is essential for the proper functioning of societies.
How can we translate research findings from social sciences gained on human-human interaction to our challenge of communications in the digital world of today and the future? In particular, how can we possibly build a model of human-digital interaction to ensure that highly interconnected digital technologies incorporate fundamental elements that enable humans to make informed trust decisions?
With 5G and 6G communications, we are at the doorstep of seeing robotic communicating devices enter our world. Lawn mowers, vacuum cleaners, and robo-cars are just the very first examples of the Tactile Internet [6]. We need to understand how we can have these robotic devices enter our lives not as untrustworthy oddballs but as trusted elements of our daily physical lives. We will then depend on the trustworthiness of the underlying communication system to build trust for our decision makings. As engineers and computer scientists, we know that successful attacks on digital systems are realistic from a technology perspective.
To comprehend the research challenge of addressing trustworthiness, we need a model that incorporates all elements and characteristics involved. We, therefore, need to translate knowledge from human-human to human-digital interactions. Using the well-known model of trustworthiness and trust from social sciences as a basis [13], this paper proposes a new model as one foundation for building trust and trustworthiness in human-digital interactions. It should serve as a basis for future research on making information and communication technologies and products truly trustworthy.
@@
In the next section, we first review the findings of social sciences. Then we look at the technical aspects of trustworthiness and their characteristics, followed by a section on the need to be able to measure. This is the basis for presenting our main contribution, i.e., a new model for trust and trustworthiness in the context of human-digital interactions and communications. Although (in this early stage of research) we have not conducted our own empirical evaluation, we provide examples from existing literature to underpin our model.
II A Brief Review of a Social Sciences Model of Trust
Figure 1: The ABI model of trustworthiness and trust by Mayer et al. [13]
@@ II A Brief Review of a Social Sciences Model of Trust
Over the last few decades, the social sciences and humanities have produced a multitude of definitions and explanations as well as empirical findings on the topic of trust [8]. Hence, we reflect on the insights from the humanities and social sciences in the area of trust and trustworthiness before conceptualizing them for the area of communications engineering.
@@
On a micro level, both psychology and educational science predominantly explore interpersonal trust in specific interaction situations. However, this micro perspective is not sufficient to convey the multiple facets of trust. A macro-level approach takes place primarily in sociology and political science, with a particular focus on system trust. Political science studies the concept of trust mainly from a developmental dynamic perspective, aiming to explain the development of trust in democracy or institutions. Economic science studies the concept of trust both from a basic theory and a problem-oriented perspective, the latter primarily in organizational and marketing research. Specifically, organizational research focuses on questions of what organizational trust is, what factors influence it, and what effects it can have on the organization. By contrast, philosophy investigates the connection between trust and morality as well as the relevance of trust for society. In comparison, communication science examines less the question of trust, but rather the credibility of the media.
@@
The spectrum of social science research on trust is very broad. Still, there has been a lack of consolidation of the research achievements on the topic of trust. Therefore, and certainly owing to the different scientific perspectives and their varying emphases, there is no uniform understanding of the term of trust. Yet, the diverse disciplinary interest shows that trust is not a simple concept, but a heterogeneous and complex subject.
@@
To investigate the concept of trust, we use the widely received integrative ABI trust model of Mayer et al. [13] on organizational trust. The focus of this model (Fig. 1), which analyzes the circumstances of trust emergence, is primarily on the trustworthiness of the trusted party as perceived by the trust giver. The authors define trust as “the willingness of a party to be vulnerable to the actions of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party” [13]. Particular emphasis is placed on the moment of vulnerability and the willingness to take a certain risk. The integrative model explains the emergence of trust in a situation in which a relationship between the two parties has not yet evolved. The factors of perceived trustworthiness influence the emergence of trust. However, the expectation of trust does not yet involve any risk, because, according to the model, risk is only the result of trust. This view is in contradiction to the views of other authors, who presume a risk perception that requires trust in the first place.
@@
According to the ABI model, the assessment of trustworthiness as the basis for an act of trust is influenced by the evaluation of the following three factors: ability, integrity, and benevolence. Abilities include skills, competencies, and attributes that enable a person to solve a specific problem. Benevolence describes how much the trusted party considers the well-being of the trusting party in his or her actions. Integrity implies qualities such as honesty and reliability. Although the factors can be considered separately, they are interrelated and can influence each other.
@@
In the ABI model, trustworthiness is understood as a continuum. At the beginning of a relationship, the trust giver has little to no information about the trust taker. In particular, the assessment of benevolence is difficult. Therefore, Mayer et al. assume that the perception of integrity is key to the initial trust building. As the relationship progresses, both actors obtain more information and gain a better assessment of ability and benevolence. Thus, as the relationship progresses positively, the degree of trust may increase.
@@
Although the ABI model is the foundation of much follow-up research on trust and trustworthiness, the number of empirical studies is still limited. A recent one, which empirically supports the ABI model with respect to innovation networks, is described in [15]. Moreover, a recent meta-study [9], which reviewed more than 300 studies, supports the ABI model by extending it with more details related to the trustee, trustor and context-environmental factors.
Parallel to the model of Mayer et al., the theory of public trust was developed in communication science, which also works with different trust factors [4]. A distinction is made between subject-specific trust factors (factual competence and problem-solving competence), socio-psychological trust factors (social behavior, communication behavior [transparency, consistency], character), and socio-normative trust factors (sense of responsibility, ethical behavior). The more completely and intensively the trust factors apply, the more likely are trust gain, trust formation or trust constitution. If the factors apply only partially and/or to a lesser extent, the greater the likelihood of a reduction or loss of trust.
@@ III Trustworthiness and its Measurability from a technical perspective
If we shift the focus towards the interaction of humans with technical systems, the definition of trust as given in the previous section can basically remain the same, i.e., the willingness of a human to be vulnerable to the actions of a technical system based on the expectation that the system will perform a particular action important to the human, irrespective of the ability to completely monitor or control the technical system. But what about the trustworthiness of the technical system and its measurability?
@@ III-A Trustworthiness
The ISO/IEC standard TS 5723:2022 [11] defines trustworthiness as the “ability to meet stakeholders’ expectations in a verifiable way”. Similarly, NIST sees trustworthiness as an objective property of a system, since the “trustworthiness of a system is based on the concept of assurance” [14]. Note that other concepts like “quality” [10], “quality of service”, or “quality of experience” have a similar broad meaning, according to their definitions. Nevertheless, it makes sense to have different terms, since the wording emphasizes different aspects. For example, quality of service is often related to the fulfillment of non-functional requirements, such as latency or bandwidth. While trustworthiness is often used interchangeably with security and privacy-related aspects, it is important to stress that the fulfillment of expectations of stakeholders of a given system spans a much broader scope of requirements for the design of a truly trustworthy system.
@@ III-B Characteristics of Trustworthiness and their measurability
However, with a broader scope, even more indicators should be included, such as:
  • Benevolence (as used in the social sciences)
@@
Moreover, the wide adoption of (in this sense, “new”) technologies such as artificial intelligence (AI), machine learning (ML), or blockchains and the related distributed ledger technology (DLT) might require an extension of the set of TCs and will influence the fundamental trustworthiness assumptions with regard to a given system. The current “black box” nature of many AI/ML algorithms leads to a situation where explainability emerges as a fundamental precondition for defining AI/ML-related TCs and, hence, ultimately for trustworthiness. Sophisticated explainability could increase the confidence in the correctness of the AI/ML results in the first place; however, it is absolutely necessary when it comes to defining and measuring TCs. An example related to large language models (LLM), such as ChatGPT, would be the output of references supporting the statements made by the LLM. A completely different technology, DLT, allows a system designer to adapt the necessary trust and trustworthiness assumptions: for example, a DLT substituting a central data store shifts the trust from a single trusted entity towards a non-malicious majority of DLT providers.
In general, a TC is a specific characteristic of a system having specific objectives. A TC is measured with the help of TC measurements, while TC controls are used to implement a given TC.
The first challenge in reasoning about trustworthiness is to have a comprehensive list of TCs. The second challenge is to define the means of measurability of the TCs. For most of the TCs, there exists no generally accepted means for their measurement. For some TCs (e.g., confidentiality or availability) the measurability appears to be more tangible than for others (like privacy or benevolence). In many cases, a combination of multiple measures is necessary to obtain a comprehensive understanding of the trustworthiness at stake. Depending on the domain and the context in which trustworthiness is being evaluated, the individual weighting of the specific TCs and the means of measurement certainly vary.
@@ IV Measurability of Trustworthiness
Ideally, trustworthiness as a whole should be measurable and quantifiable to allow for comparing systems and to meet the requirement of verifiability of user expectations. There exist various proposals in the literature to create a metric for trustworthiness (e.g., [5] or the work of ITU-T on “Trustworthiness evaluation for autonomous networks including IMT-2020 and beyond”). Yet there is still no holistic trustworthiness metric that can be applied universally. The challenges behind designing a metric for trustworthiness are manifold. One reason is that trustworthiness involves several different TCs – and many of them lack sound metrics. Take, for instance, privacy as one TC. Privacy itself comprises different aspects, such as anonymity and unlinkability. However, even for these more basic properties of privacy, no universal applicable metrics exist besides isolated measures like $k$-anonymity or differential privacy [17].
@@
Another challenge is related to the intended nature of a metric for trustworthiness. Should it be a single number allowing for easy comparison but limiting expressiveness? Should it be a vector containing a value for each TC? This would enhance the expressiveness but makes comparison more difficult. Taking consumer reports as an example, we can often see a combination of both approaches: a single value (e.g., inspired by school marks) for easy and quick assessments, but, additionally, measurements for individual properties of the service or product in question, allowing a consumer to adapt the “base assessments” according to their individual situation, beliefs, and requirements. Similar approaches can also be seen in the IT security domain. Here, the Common Vulnerability Scoring System (CVSS), on the one hand, provides a single score, allowing system administrators to do a quick risk assessment. But, on the other hand, it also allows for individual adaptation based on individual requirements (e.g., which protection goals are important) and the individual situation (e.g., do countermeasures exist like firewalls, which mitigate exploitability). Would it, therefore, make sense to define a trustworthiness vector based on the measured TCs and give a weighted trustworthiness score?
@@
Another question regarding a sound trustworthiness metric is whether the trustworthiness value should be time-dependent or time-invariant. A time-dependent metric might be more intuitive in the first place since the assessment of whether, e.g., a cryptographic algorithm can be considered trustworthy might change over time when new weaknesses of the algorithm are discovered (thanks to cryptanalysis), or due to new technological developments (e.g., quantum computers). However, a time-dependent metric would be rather useless if the related values were valid only in the very moment they were measured (or calculated), much less help guide humans in their trust decisions.
@@
A trustworthiness metric needs to be designed in a way that leads to values which have a reasonable period of validity. Consequently, a metric which leads to time-independent values (or at least to values which equally change over time for similar systems) would be desired. Taking a second look at the example of the cryptographic algorithm and, specifically, the symmetric encryption algorithm DES, one could argue that the inherent security “level” provided by DES in terms of effort necessary to find the secret key (through calculations) has not changed substantially since its introduction in the 1970s. What changes is the “perceived” or relative security, given the technological advancements in computation power allowing (investing the same amount of money) to break DES today within minutes, while in the 1970s it took years. Translating this to another domain: What was considered a “fast” car in the 1930s might be considered a “slow” car today – although the maximum speed of an individual car is time-invariant.
@@
Given the observations above, a possible solution could be to have two kinds of trustworthiness metrics: one “absolute” trustworthiness value, which is time-invariant to reflect the essence of trustworthiness provided by the system and to allow for comparison, and a “relative” trustworthiness value, which is time-dependent but captures better the current level of trustworthiness to more easily support human trust decisions. Besides, the need for the latter can even be derived from the definition of trustworthiness itself, which is the “ability to meet stakeholders’ expectations”. Therefore, the “level of trustworthiness” can always be understood as relative to the expectation of a given user (stakeholder).
@@
By reviewing the list of the different TCs in the previous section, it becomes apparent that the TCs are not necessarily independent of each other. This poses additional challenges in optimizing the overall trustworthiness of a system. Consider, for example, TCs which are directly related to protection goals from the IT security domain, e.g., confidentiality, integrity, accountability, availability, anonymity, etc. There are certain dependencies among these protection goals [19]. For example, confidentiality and availability weaken each other. Another example: Given the current state of the art, security enhancements often have a negative impact on usability; increased privacy could lead to reduced functionality expressed by the “privacy vs. utility” trade-off during system design. Consequently, for an informed trust decision that we can rely on, the correlations of the TCs must be analyzed and taken into account. It is important to understand specific definitions and measuring methods of the TCs to ensure the filtering and weighting of the TCs result in a meaningful trust decision that confirms the learnings and minimizes risks.